OpenStudy (unklerhaukus):

\[2^4+10^3+3^3\times37\]

hartnn (hartnn):

37*9 is special adding 1032 to tht gives 2015 :)

hartnn (hartnn):

37*3 37*6 37*9 ... a special pattern ;)

hartnn (hartnn):

**adding 1016 :P

OpenStudy (unklerhaukus):

\[2^4+10^3+3^3\times37\\16+1000+3^2\times3\times37\\1016+9\times111\\1016+999\\\]

hartnn (hartnn):

\(\Large 37\times 3 =111 \) \(\Large 37\times 6 =222 \) \(\Large 37\times 9 =333 \) \(\Large 37\times 12 =444 \) \(\Large 37\times 15=555 \) \(\Large 37\times 18 =666 \) \(\Large 37\times 21 =777 \) \(\Large 37\times 24 =888 \) \(\Large 37\times 27 =99 \)

hartnn (hartnn):

\(\Large 37\times 27 =999\) ***

OpenStudy (unklerhaukus):

What is an expression for 2015 in only prime numbers

OpenStudy (unklerhaukus):

Something better than \[2^4+2^35^3+3^3\times37\]

ganeshie8 (ganeshie8):

2015 is a square free integer

hartnn (hartnn):

\(\Large 5\times 13 \times 31\)

OpenStudy (anonymous):

is there a special number have same property as 37 ?

OpenStudy (anonymous):

makes me remember this 37 373737 3737373737 37373737373737 .....

OpenStudy (anonymous):

all composite

OpenStudy (anonymous):

this is the rule (why its composite) 37(10101...).

hartnn (hartnn):

a 3 digit number , say pqr. \(\Large pqr \times 1001 = pqrpqr\)

ganeshie8 (ganeshie8):

\[\begin{vmatrix}5 & 2 & 7 \\ 0 & 13 & 11 \\ 0 & 0 & 31\end{vmatrix} = 2015\]

hartnn (hartnn):

\(\begin{vmatrix}5 & \heartsuit & \clubsuit \\ 0 & 13 & \spadesuit \\ 0 & 0 & 31\end{vmatrix} = 2015\) :P

OpenStudy (unklerhaukus):

neat

sammixboo (sammixboo):

2014 + 1 = 2015 :D

hartnn (hartnn):

omg, why didn't i think of that! :P smart

sammixboo (sammixboo):

Ik ik ~*bows*~

ganeshie8 (ganeshie8):

Clever ;) by goldbach conjecture it can be represented as sum of 3 primes

OpenStudy (unklerhaukus):

What is the sum of primes that equals 2015?

sammixboo (sammixboo):

Nope

sammixboo (sammixboo):

659 + 673 + 683 ?

sammixboo (sammixboo):

No, dang it. Lol I am trying

OpenStudy (unklerhaukus):

2011+2+2

sammixboo (sammixboo):

Oh my goodness I feel like somebody just smacked me with a bat

OpenStudy (unklerhaukus):

659 + 673 + 683 is way better.

OpenStudy (anonymous):

:|

sammixboo (sammixboo):

661 + 643 + 3*2^3 + 673 + 7*2 I am getting bored as you can tell. I just created a few equations on how to get 2015. I am working on a more complex one right now

ganeshie8 (ganeshie8):

hey no wait there are like thousands of ways in which you can represent 2015 as sum of 3 different primes.. here is a more fun one : represent 2015 as sum of squares of two integers

ganeshie8 (ganeshie8):

\[2015 = a^2 + b^2\] find \(a\) andn \(b\)

sammixboo (sammixboo):

I don't know but this is the closest I have gotten xD 35.2^2 + 27.87^2

sammixboo (sammixboo):

To 2015. I am just guessing right now. I will probably try to work it out when I am fully awake

OpenStudy (anonymous):

eh no integer solution (face palm)

OpenStudy (anonymous):

any way this circle gives all solution :P 

ganeshie8 (ganeshie8):

lol does that mean all the points on that circle have irrational coordinates ? it turns out that 2015 is a special number which cannot be represented as sum of squares of two integers and the reason for that goes back to 17th century https://www.math.hmc.edu/funfacts/ffiles/20008.5.shtml

OpenStudy (anonymous):

infact i check from 1 to 45

OpenStudy (anonymous):

checked*

ganeshie8 (ganeshie8):

bruteforce is for programmers, not mathematicians :P

OpenStudy (anonymous):

who cares

OpenStudy (anonymous):

but yes i likes 4k+3 and 4k+1 idea since 2015 is 3 mod 4 i completely forgot about that

OpenStudy (anonymous):

but u seems like exited about 2015 :P @ganeshie8 note that its a happy number

ganeshie8 (ganeshie8):

2015 -> 2^2 + 0^2 + 1^2 + 5^2 = 4+0+1+25 = 30 -> 3^2 + 0^2 = 9 -> 9^2 = 81 -> 8^2 + 1^2 = 64+1 = 65 -> 6^2 + 5^2 = 36 + 25 = 61 -> 6^2+1^2 = 37 im tired :O

OpenStudy (anonymous):

well 3 is a happy number i stopped when i got 30 :P

ganeshie8 (ganeshie8):

how do u know 3 is happy

OpenStudy (anonymous):

but lets continue 37=9+49=58 58=25+64=89 89=64+81=145 145=1+16+25=42 42=16+4=20 20=4 4=16=1+36=37 seems like sad not happy

ganeshie8 (ganeshie8):

does that mean 3 is also not happy ? because all the numbers in the sequence of steps have to fall on same branch

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i would say a happy number can be written as sum of 2 squares

ganeshie8 (ganeshie8):

counter example : 19

OpenStudy (anonymous):

hmm or sum of squares can be a happy number :P